OpenStudy (anonymous):

A baseball player hits a baseball into the outfield. The following equation models the ball's path. If not caught, when will the baseball land? Round to the nearest tenth second. h = -16t2 + 50t + 4 A. 3.0 seconds B. 3.2 seconds C. 3.4 seconds D. 3.6 seconds

OpenStudy (anonymous):

h=0 when the baseball lands -16t^2+50t+4=0 -8t^2+25t+2=0 use the quadratic formula and discard the negative answer

OpenStudy (anonymous):

B. 3.2 seconds